Define r.U and Mathematical expectation of a r.U. A discrete r.U. has the following probability distribution.
X : -2 -1 0 1 2 3
P(X) : 0.1 a 0.2 2a 0.3 a
Find i) The value of ‘a’,
ii) its distribution function F(X),
iii) Mean and variance of X.
1
Expert's answer
2020-08-17T18:03:30-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 x & -2 & -1 &\ \ 0\ & \ 1\ & 2\ & \ 3 \\ \hline
 p(x) & 0.1 & a & 0.2 & 2a & 0.3 \ & a
\end{array}"

(i)

"\displaystyle\sum_{x}p(x)=1""0.1+a+0.2+2a+0.3+a=1""4a=0.4""a=0.1"

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c}
 x & -2 & -1 &\ \ 0\ & \ 1\ & 2\ & \ 3 \\ \hline
 p(x) & 0.1 & 0.1 & 0.2 & 0.2 & 0.3 \ & 0.1
\end{array}"

(ii)

"F(-2)=F(X\leq-2)=p(-2)=0.1"


"F(-1)=F(X\leq-1)=p(-2)+p(-1)="

"=0.1+0.1=0.2"

"F(0)=F(X\leq0)=p(-2)+p(-1)+p(0)="

"=0.1+0.1+0.2=0.4"


"F(1)=F(X\leq1)=p(-2)+p(-1)+p(0)+p(1)="

"=0.1+0.1+0.2+0.2=0.6"


"F(2)=F(X\leq2)=p(-2)+p(-1)+p(0)+p(1)+p(2)="

"=0.1+0.1+0.2+0.2+0.3=0.9"


"F(3)=F(X\leq3)=1"


"F(x) = \begin{cases}
 0 & \ \ \ x<-2 \\
 0.1 & -2\leq x<-1 \\
 0.2 & -1\leq x<0 \\
 0.4 &\ \ \ 0\leq x<1\\
 0.6 & \ \ \ 1\leq x<2 \\
 0.9 & \ \ \ 2\leq x<3\\
 1 &\ \ \ 3\leq x 
\end{cases}"

(iii)


"\mu=E(X)=\displaystyle\sum_{x}x\cdot p(x)"

"\mu=E(X)=-2\cdot0.1+(-1)\cdot0.1+0\cdot0.2+""+1\cdot0.2+2\cdot0.3+3\cdot0.1=0.8"

"\sigma^2=E(X^2)-\mu^2"

"E(X^2)=(-2)^2\cdot0.1+(-1)^2\cdot0.1+(0)^2\cdot0.2+""+(1)^2\cdot0.2+(2)^2\cdot0.3+(3)^2\cdot0.1=2.8"

"\sigma^2=2.8-(0.8)^2=2.16"